OpenStudy (anonymous):

Solve 5/2cos2x-1/2=3sinx to two decimal places where x is equal or between 0 and 2pi

OpenStudy (p0sitr0n):

expand cos2x and pair it up with sinx to get tanx and then check where the solution is located on the function

OpenStudy (anonymous):

can you rewrite your question in equation, so it will be more clear?

OpenStudy (anonymous):

the equation is \[\frac{ 5 }{ 2 }\cos^2x-\frac{ 1 }{ 2 }=3sinx\]

OpenStudy (anonymous):

@mihirb ok

OpenStudy (solomonzelman):

Oh, have you ever heard of this? \[sinx=\sqrt{1-Cos^2x}~~~~~~~~~~and\]\[cosx=\sqrt{1-Sin^2x}\]

OpenStudy (solomonzelman):

Good tip no?

OpenStudy (solomonzelman):

\[5Cos^2x-1=6Sinx\]\[5(\sqrt{1-Sin^2x})-1=6Sinx~~~~~~~~l et~~Sinx=a\]

OpenStudy (solomonzelman):

\[5\sqrt{1-a^2}-1=6a,~~~~~add~~~1~~~t o~~~both~~~sides,\]\[5\sqrt{1-a^2}=6a+1~~~~~~~squa re~~~both~~~sides,\]\[25(1-a^2)=36a^2+12a+1~~~~~~->~~~~~25-25a^2=36a^2+12a+1~~~~->\]\[61a^2+12a+26~~~~solve~~for~~a~~and~~substitute~~the~~result~~into~~Sinx=a\]